Assume that 0.271 g of diborane is combusted in a calorimeter whose heat capacity (Ccalorimeter) is 7.854 kJ/°C at 23.60°C. What is the final temperature of the calorimeter?
1
Expert's answer
2020-12-03T13:53:50-0500

Combustion of diborane is occur according to the following equation,


B2H6 + 3O2 ----> B2O3 + 3H2O

Heat of combustion, ∆H = –2035 kJ/mol

(–) is due to liberation of heat


The heat absorbs by the calorimeter,

                            H = 2035 kJ

Heat capacity of calorimeter,S = 7.854 kJ/°C

Initial temperature of calorimeter,

                            T1 = 23.60°C

Let, final temperature of calorimeter = T2°C


So,  H = S×(T2–T1)

 Or, (T2–T1) = H/S

 Or, T2 = (H/S) + T1

       = [(2035/7.854)+23.70]

       = 282.7 ° C

       

Hence, find temperature of calorimeter is 282.7 °C.